# The Gradient of Mean Molecular Weight Across the Radius Valley

### v1

@@ Abstract
Photo-evaporation shapes the observed radii of small exoplanets and constrains the underlying distributions of atmospheric and core masses. However, the diversity of atmospheric chemistries corresponding to these distributions remains unelucidated. We develop a first-principles carbon-hydrogen-oxygen-sulfur-silicon (CHOSSi) outgassing model that accounts for non-ideal gas behavior (via fugacities) at high pressures, as well as the tendency for water and hydrogen to dissolve in melt (via solubility laws). We use data-driven radius valley constraints to establish the relationship between the atmospheric surface pressures and melt temperatures of sub-Neptunes. Sub-Neptunes with less massive rocky cores retain less of their primordial hydrogen envelopes, which leads to less heat retention and diminished melt temperatures at the surfaces of these cores. Lower melt temperatures lead thermodynamically to the dominance of carbon-, oxygen-, sulfur- and silicon-bearing molecules over molecular hydrogen, which naturally produce a diversity of mean molecular weights. Our geochemical outgassing calculations robustly predict a gradient of mean molecular weight across the radius valley, where the strength of this gradient is primarily driven by the oxygen fugacity of the molten cores and not by the carbon enrichment (or “metallicity”) of the atmosphere. Smaller sub-Neptunes are predicted to have less hydrogen-dominated atmospheres. The precise relationship between the observed and outgassed chemistries requires an understanding of how convection near the core interacts with large-scale atmospheric circulation (driven by stellar heating) near the photosphere, as well as the influence of photochemistry.
@@ 1 Introduction
With mean molecular weights of intermediate values being reported for small exoplanets (e.g., $\mu=5.47^{+1.25}_{-1.14}$ for the sub-Neptune TOI-270d; Benneke et al. 2024), it becomes relevant to understand if hybrid atmospheres may account for them. As the primordial hydrogen- helium envelope becomes more massive, we expect the mean molecular weight to tend towards $\mu\approx 2.2$.
@@ 2.2.1 Solubility laws
Outgassing models typically describe a system consisting of a mixed gas (the atmosphere) and a liquid (the melt) in collective chemical equilibrium (e.g., French 1966; Gaillard & Scaillet 2014; Gaillard et al. 2022; Tian & Heng 2024). Some of these gases are partially dissolved in the melt, a phenomenon that is well established in the geosciences (e.g., Gaillard et al. 2022). In the current study, we take the approximation that only water (H2O) and molecular hydrogen (H2) have non-negligible solubilities. The former is well established in the geosciences literature (e.g., McMillan 1994; Berndt et al. 2002; Dufils et al. 2020). The latter is motivated by the claim that objects with sizes larger than about three Earth radii are rare because of the solubility of molecular hydrogen in melt under high pressure—the so-called “fugacity crisis” (Kite et al., 2019). Under Earth-like conditions, carbon dioxide (CO2), carbon monoxide (CO) and methane (CH4) have lower solubilities, compared to water, in basaltic melts (e.g., Amalberti et al. 2021). We also ignore the solubility of sulfur species as the solubility laws are unavailable.
@@
Figure 2 illustrates the importance of including non-unity fugacity coefficients ($\phi_{i}\neq 1$) in the solubility laws of both water and hydrogen. At 1600 K, the discrepancy for the mass fraction of hydrogen dissolved in the melt may be an order of magnitude if one sets $\phi_{i}=1$. Generally, the mass fraction of hydrogen dissolved in the melt is small: $\sim 1\%$ at $\sim 10$ kbar. The mass fraction of water dissolved in melt depends on its mixing ratio, but is generally well below $1\%$ unless the atmosphere becomes water-dominated. As expected, the solubility depends predominantly on pressure and has a weak dependence on temperature.
@@ 2.2.2 Net chemical reactions and equilibrium constants
$\begin{split}\mbox{CO}+\frac{1}{2}\mbox{O}_{2}&\Longleftrightarrow\mbox{CO}_{2},\\ \mbox{H}_{2}+\frac{1}{2}\mbox{O}_{2}&\Longleftrightarrow\mbox{H}_{2}\mbox{O},\\ \mbox{CH}_{4}+2\mbox{O}_{2}&\Longleftrightarrow\mbox{CO}_{2}+2\mbox{H}_{2}\mbox{O},\\ \frac{1}{2}\mbox{S}_{2}+\mbox{O}_{2}&\Longleftrightarrow\mbox{SO}_{2},\\ \mbox{H}_{2}\mbox{S}+\frac{1}{2}\mbox{O}_{2}&\Longleftrightarrow\frac{1}{2}\mbox{S}_{2}+\mbox{H}_{2}\mbox{O}.\\ \end{split}$ (5)
@@
$\begin{split}\mbox{CH}_{4}+2\mbox{H}_{2}\mbox{O}&\Longleftrightarrow\mbox{CO}_{2}+4\mbox{H}_{2},\\ \mbox{CO}+\mbox{H}_{2}\mbox{O}&\Longleftrightarrow\mbox{H}_{2}+\mbox{CO}_{2},\\ \mbox{CH}_{4}+\mbox{H}_{2}\mbox{O}&\Longleftrightarrow\mbox{CO}+3\mbox{H}_{2},\\ \end{split}$ (6)
@@
$\mbox{SiO}+3\mbox{H}_{2}\Longleftrightarrow\mbox{SiH}_{4}+\mbox{H}_{2}\mbox{O}.\$ (7)
@@ 2.2.3 Thermodynamic quantities
where the index $j$ runs from 1 to 6 and ${\cal R}$ is the universal gas constant. In Appendix A, we describe how the Gibbs free energies are obtained and provide fitting functions for $\Delta G_{j}$ corresponding to the 6 net chemical reactions in equations (5) and (7). Details for how to construct $\Delta G$ for net reactions have previously been provided (e.g., Heng & Lyons 2016).
@@ 2.2.5 Reconciling Dalton’s law with Newton’s second law
where $M_{\rm atm}$ is the mass of the atmosphere, $g$ is the surface gravity of the exoplanet and $R_{\rm core}$ is its radius of the core. Let $M_{i}$ be the mass of species $i$ present in the atmosphere. By analogy, it is tempting to write a similar expression for the partial pressure, but this comes with the implication that
@@
To proceed, we need to clarify a potential source of confusion concerning terminology across different scientific disciplines. The molecular mass, defined in this study as $m_{i}$, is simply the mass of a molecule of species $i$. It has physical units of mass. If we write the atomic mass unit (amu) as $m_{\rm u}$, then the molecular weight is given by $\mu_{i}=m_{i}/m_{\rm u}$. It is a dimensionless quantity. A third quantity, the so-called “molar mass” (used in, e.g., Bower et al. 2019), is the average mass of an ensemble of particles of species $i$ and has physical units of mass per mole (g mol-1). All three quantities have their counterparts when averaged over the entire chemical system (containing an arbitrary number of species): mean molecular mass ($\bar{m}$), mean molecular weight ($\mu$) and mean molar mass. In the current study, we utilize only $m_{i}$, $\bar{m}$, $\mu_{i}$ and $\mu$ to minimize confusion.
@@ 2.2.6 Hydrogen budget
$\begin{split}A^{\prime}_{\rm H_{2}O}&\equiv A_{\rm H_{2}O}\left(\frac{\phi_{\rm H_{2}O}}{P_{0}}\right)^{1/2},\\ A^{\prime}_{\rm H_{2}}&\equiv\frac{A_{\rm H_{2}}\phi_{\rm H_{2}}}{P_{0}}.\\ \end{split}$ (23)
@@ 2.3 Solution method
[…]
@@ 3.2 Model design and choice of parameters
One of the lessons learned in the current study is the relevance of model design: how the equations, unknowns and parameters are defined. There are an arbitrary number of ways to formulate this. Instead of parametrizing the bulk carbon, hydrogen, carbon and silicon content of an exoplanet (which is generally unknown), we describe it by ratios of elemental abundances representing either the atmosphere or the melt (that outgasses the atmosphere). We mainly focus on hybrid atmospheres and designate H2O, CO, CO2, CH4, H2S, SO2, SiO and SiH4 as the unknown gaseous species whose abundances that we wish to solve for. Since there are 5 net chemical reactions, an equation for C/H, an equation for Si/O, the condition that the partial pressures must sum up to the total pressure and the expression for the mean molecular weight, this formally constitutes 9 equations and 9 unknowns.
@@
    Oxygen fugacity ($f_{\rm O_{2}}$): The oxygen fugacity is the effective partial pressure of oxygen gas participating in equilibrium reactions involving multi-valent elements like iron and silicon in the melt. In other words, the oxygen fugacity is buffered by the melt. It is typically specified relative to a chemical buffer, because the amount of gaseous oxygen liberated depends on temperature. A commonly used buffer is that of iron-wüstite (IW) (Wade & Wood, 2005),
@@
    The upper mantle of the Earth is estimated to have an oxygen fugacity of $\log{f_{\rm O_{2}}}=\mbox{IW}+3.5$ (Frost & McCammon, 2008), which corresponds to $f_{\rm O_{2}}\sim 10^{-7}$ bar for $T=1600$ K and $P=1$ bar.
    The approach of specifying the oxygen fugacity relative to the IW buffer cannot be implemented using existing empirical formulae, because they are not calibrated for temperatures well above 2000 K. As a striking example, if IW$+3.5$ is extrapolated to 5000 K, using the empirical relationship of Ballhaus et al. (1991) as stated in equation (27) of Tian & Heng (2024), then it corresponds to $f_{\rm O_{2}}\approx 25$ kbar for $P=10$ kbar, i.e., the amount of gaseous oxygen liberated exceeds the atmospheric surface pressure, which is physically nonsensical.
@@
[…]
    The Solar System gives no guidance on how to predict the oxygen fugacity of a rocky body with a mass exceeding that of Earth. If we use the estimated values for the upper mantles of Earth (IW$+3.5$) and Mars (IW) (Deng et al., 2020), then a crude, empirical scaling relationship may be obtained,
@@
    For a rocky core of 4 Earth masses, this scaling relationship produces an implausible value of IW$+15.3$, which corresponds to about $70$ kbar at 1600 K.
@@
[…]
@@
[…]
@@
Figure 7: Outgassing calculations informed by data-driven radius valley constraints. Left panel: reduced conditions corresponding to $\log{f_{\rm O_{2}}}$ and $\log{f_{\rm S_{2}}}$ being uniformly sampled between $-11$ to $-1$ (cgs units; corresponding to $10^{-17}$ to $10^{-7}$ bar). Right panel: oxidized conditions corresponding to $\log{f_{\rm O_{2}}}$ and $\log{f_{\rm S_{2}}}$ being uniformly sampled between $-1$ to $2$ (cgs units; corresponding to $10^{-7}$ to $10^{-4}$ bar). The elemental abundance of carbon is assumed to be solar (see text for more details).
@@ 3.3 Representative case studies of hybrid atmospheres
[…]
@@
There is a trend of the mean molecular weight increasing from $\mu\approx 2$ to $\mu\approx 20$ as the melt temperature decreases. This is slightly more pronounced for Earth-like, compared to the sub-Neptune-like, surface pressures. Nevertheless, the dominant effect is the diminished temperature, rather than the diminished pressure. Physically, these two properties are closely related: as the core becomes less massive, it retains less of its hydrogen envelope due to weaker gravity. This in turn allows the exoplanet to cool more quickly.

### v2

@@ Abstract
Photo-evaporation shapes the observed radii of small exoplanets and constrains the underlying distributions of atmospheric and core masses. However, the diversity of atmospheric chemistries corresponding to these distributions remains unelucidated. We develop a first-principles carbon-hydrogen-oxygen-sulfur-silicon (CHOSSi) outgassing model that accounts for non-ideal gas behavior (via fugacities) at high pressures, as well as the tendency for water and hydrogen to dissolve in melt (via solubility laws). We use data-driven radius valley constraints to establish the relationship between the atmospheric surface pressures and melt temperatures of sub-Neptunes. Sub-Neptunes with less massive rocky cores retain less of their primordial hydrogen envelopes, which leads to less heat retention and diminished melt temperatures at the surfaces of these cores. Lower melt temperatures lead thermodynamically to the dominance of carbon-, oxygen-, sulfur- and silicon-bearing molecules over molecular hydrogen, which naturally produce a diversity of mean molecular weights. Our geochemical outgassing calculations robustly predict a gradient of mean molecular weight across the radius valley, where the strength of this gradient is primarily driven by the oxygen fugacity of the molten cores and not by the carbon enrichment (or “metallicity”) of the atmosphere. Smaller sub-Neptunes are predicted to have less hydrogen-dominated atmospheres. Establishing the precise relationship between the observed and outgassed chemistries requires an understanding of how convection near the core interacts with large-scale atmospheric circulation (driven by stellar heating) near the photosphere, as well as the influence of photochemistry.
@@ 1 Introduction
With mean molecular weights of intermediate values being reported for small exoplanets (e.g., $\mu\approx 5$66 for the sub-Neptune TOI-270d; Benneke et al. 2024; Felix et al. 2025), it becomes relevant to understand if hybrid atmospheres may account for them. As the primordial hydrogen- helium envelope becomes more massive, we expect the mean molecular weight to tend towards $\mu\approx 2.2$.
@@ 2.2.1 Solubility laws
Outgassing models typically describe a system consisting of a mixed gas (the atmosphere) and a liquid (the melt) in collective chemical equilibrium (e.g., French 1966; Gaillard & Scaillet 2014; Gaillard et al. 2022; Tian & Heng 2024). Some of these gases are partially dissolved in the melt, a phenomenon that is well established in the geosciences (e.g., Gaillard et al. 2022). In the current study, we take the approximation that only water (H2O) and molecular hydrogen (H2) have non-negligible solubilities. The former is well established in the geosciences literature (e.g., McMillan 1994; Berndt et al. 2002; Dufils et al. 2020). The latter is motivated by the claim that objects with sizes larger than about three Earth radii are rare because of the dissolution of molecular hydrogen in melt under high pressure—the so-called “fugacity crisis” (Kite et al., 2019). Under Earth-like conditions, carbon dioxide (CO2), carbon monoxide (CO) and methane (CH4) have lower solubilities, compared to water, in basaltic melts (e.g., Amalberti et al. 2021; Seo et al. 2024). Section 2.3 of Seo et al. (2024) contains a comprehensive discussion of the solubility laws associated with CO2, CO and CH4. Using scaling arguments, Seo et al. (2024) conclude that the dissolution of CO and CH4 into the melt is negligible, while CO2 is moderately soluble when the atmosphere is “highly oxidized” and the silicate layer of the core is “almost fully molten and well-mixed”. We also ignore the solubility of sulfur species as the solubility laws are unavailable.
@@
Figure 2 illustrates the relevance of including non-unity fugacity coefficients ($\phi_{i}\neq 1$) in the solubility laws of both water and hydrogen. At 1600 K, the discrepancy for the mass fraction of hydrogen dissolved in the melt may be an order of magnitude if one sets $\phi_{i}=1$. Generally, the mass fraction of hydrogen dissolved in the melt is small: $\sim 1\%$ at $\sim 10$ kbar. The mass fraction of water dissolved in melt depends on its mixing ratio, but is generally well below $1\%$ unless the atmosphere becomes water-dominated. As expected, the solubility depends predominantly on pressure and has a weak dependence on temperature.
@@ 2.2.2 Net chemical reactions and equilibrium constants
[…]
@@
[…]
@@
$\mbox{SiO}+3\mbox{H}_{2}\Longleftrightarrow\mbox{SiH}_{4}+\mbox{H}_{2}\mbox{O}.\\$ (7)
@@ 2.2.3 Thermodynamic quantities
[…]
@@ 2.2.5 Reconciling Dalton’s law with Newton’s second law
where $M_{\rm atm}$ is the mass of the atmosphere, $g$ is the surface gravity of the exoplanet and $R_{\rm core}$ is its radius of the core [^5] Note that this refers to the entire rocky core of the exoplanet and not just the metallic core as is the convention of the Earth sciences.. Let $M_{i}$ be the mass of species $i$ present in the atmosphere. By analogy, it is tempting to write a similar expression for the partial pressure, but this comes with the implication that
@@
To proceed, we need to clarify a potential source of confusion concerning terminology across different scientific disciplines. The molecular mass, defined in this study as $m_{i}$, is simply the mass of a molecule of species $i$. It has physical units of mass. If we write the atomic mass unit (amu) as $m_{\rm u}$, then the molecular weight is given by $\mu_{i}=m_{i}/m_{\rm u}$. It is a dimensionless quantity. A third quantity, the so-called “molar mass” (used in, e.g., Bower et al. 2019), is the average mass of an ensemble of particles of species $i$ and has physical units of mass per mole (g mol-1). All three quantities have their counterparts when averaged over the entire chemical system (containing an arbitrary number of species): mean molecular mass ($\bar{m}$), mean molecular weight ($\mu$) and mean molar mass. In the current study, we utilize only $_{i}, m¯\bar{m}$, $\mu_{i}$ and $\mu$ to minimize confusion.
@@ 2.2.6 Hydrogen budget
[…]
@@ 2.2.7 Carbon, oxygen and silicon budgets
It is worth noting that $P_{\rm O}$ and $P_{\rm Si}$ include only silicon and oxygen in the atmosphere (gas phase), an assumption that will be justified in Section 3.2.
@@ 2.3 Solution method
where $\mbox{C/H}\equiv N_{\rm C}/N_{\rm H}$ is the elemental abundance of carbon (relative to hydrogen). Its solar value is $\mbox{C/H}=2.5\times 10^{-4}$. The ratio of silicon to oxygen (by number) is given by $\mbox{Si/O}\equiv N_{\rm Si}/N_{\rm O}$. For example, Misener et al. (2023) considered the conversion of gaseous silica (liquid SiO2) to silicon monoxide and assumed $\mbox{Si/O}=0.5$ for their chemical system.
@@ 3.2 Model design and choice of parameters
One of the lessons learned in the current study is the relevance of model design: how the equations, unknowns and parameters are defined. There are an arbitrary number of ways to formulate this. Instead of parametrizing the bulk carbon, hydrogen, carbon and silicon content of an exoplanet (which is generally unknown), we describe it by ratios of elemental abundances representing either the atmosphere or the melt (that outgasses the atmosphere). We focus exclusively on hybrid atmospheres and designate H2O, CO, CO2, CH4, H2S, SO2, SiO and SiH4 as the unknown gaseous species whose abundances we wish to solve for. Since there are 5 net chemical reactions, an equation for C/H, an equation for Si/O, the condition that the partial pressures must sum up to the total pressure and the expression for the mean molecular weight, this formally constitutes 9 equations and 9 unknowns.
@@
    Oxygen fugacity ($f_{\rm O_{2}}$): The oxygen fugacity is the effective partial pressure of oxygen gas participating in equilibrium reactions involving multi-valent elements like iron in the melt. In other words, the oxygen fugacity is buffered by the melt. It is typically specified relative to a chemical buffer, because the amount of gaseous oxygen liberated depends on temperature. A commonly used buffer is that of iron-wüstite (IW) (Wade & Wood, 2005),
@@
    The upper mantle of the Earth is estimated to have an oxygen fugacity of $\log{f_{\rm O_{2}}}=\mbox{IW}+3.5$ (Frost & McCammon, 2008), which corresponds to $f_{\rm O_{2}}\sim 10^{-7}$ bar for $T=1600$ K and $P=1$ bar (Ballhaus et al., 1991; Hirschmann, 2021).
    Strictly speaking, the approach of specifying the oxygen fugacity relative to the IW buffer cannot be implemented using existing empirical formulae, because they are not calibrated for temperatures above 3000 K (Ballhaus et al., 1991; Hirschmann, 2021) (see Appendix D). As a striking example, if IW+3.5+3.5 is extrapolated to 5000 K using the empirical formula of Hirschmann (2021), then it corresponds to $f_{\rm O_{2}}\approx 28.5$ kbar for a pressure of $10$ kbar, i.e., the amount of gaseous oxygen liberated exceeds the atmospheric surface pressure.
@@
    Figure 6: Outgassing calculations for sub-Neptune-like ($P_{\rm H_{2}}=10$ kbar; top row) and Earth-like ($P_{\rm H_{2}}=1$ bar; bottom row) surface pressures. For illustration, we have assumed the oxygen and sulfur fugacities to be equal and have the value $f_{\rm O_{2}}=f_{\rm S_{2}}=10^{-1}$ dyne cm-2 $10^{-7}$ bar. The elemental carbon abundance is set to its solar value ($\mbox{C/H}=2.5\times 10^{-4}$), while the silicon-to-oxygen ratio is set to $\mbox{Si/O}=0.3$. The melt mass fraction has been set to $\chi_{\rm melt}=100$. The dotted curves are additional calculations that assume ideal gas behavior ($\phi_{i}=1$) wherever possible (for H2, H2O, CO, CO2 and CH4), while the dashed curves ignore the solubility laws (for H2 and H2O). Left column: volume mixing ratios of various molecules as functions of melt temperature. Non-ideal gas behavior is only relevant for the sub-Neptune-like surface pressure. Right column: mean molecular weight as a function of melt temperature. For the sub-Neptune-like case study, the $\log{f_{\rm O_{2}}}-\mbox{IW}$ values range from 8.7-8.7 to 0.2-0.2 across the temperature range of 2000200060006000 K. For the Earth-like case study, it ranges from 8.5-8.5 to $0.1$.
    The Solar System gives no guidance on how to predict the oxygen fugacity of a rocky body with a mass exceeding that of Earth. If we use the estimated values for the upper mantles of Earth (IW+3.5+3.5) and Mars (IW) (Deng et al., 2020), then a crude, empirical scaling relationship may be obtained,
@@
    For a rocky core of 4 Earth masses, this scaling relationship produces an implausible value of IW+15.3+15.3, which corresponds to about $70$ kbar at 1600 K.
    A predictive theory for the formation of sub-Neptunes would elucidate how $\log{f_{\rm O_{2}}}-\mbox{IW}$ behaves as a function of the age, mass, radius, orbital period and the other properties of a sub-Neptune—without hiding our ignorance behind dozens of free parameters. A survey of the literature reveals that such a predictive theory remains elusive. For example, Lichtenberg (2021) sketches a qualitative scheme for describing the oxidation state of the mantle, but uses order-of-magnitude scaling relations, which lack predictive power, to describe the unsolved problem of atmospheric mixing (see Section 4.4.4). As another example, Kite et al. (2020) considers end members for the core and magma ocean compositions of sub-Neptunes. In their Section 4.2, Kite et al. (2020) ruminates: “Sub-Neptune formation is not well understood. We do not know how rocky cores form”; “We do not know where the growth happened”; “We do not know where the volatiles came from—nebula gas or solid-derived volatiles?”
@@
    Elemental carbon abundance (C/H): Carbon tends to reside in the gaseous phase rather than be dissolved in the melt. To lowest order, the carbon-to-hydrogen ratio (by number) describes the enrichment of the primordial hydrogen envelope. Its solar-abundance value is $\mbox{C/H}=2.5\times 10^{-4}$, which we use as a plausible starting point.
@@
    Silicon-to-oxygen ratio of melt (Si/O): The atmosphere is assumed to have the same Si/O ratio as the melt. If a pure silica (SiO2) melt is assumed, then one obtains $\mbox{Si/O}=0.5$ (Misener et al., 2023). A more realistic basaltic melt contains oxides of silicon, iron, magnesium, etc, as well as dissolved water (cf. Table 1 on pg. 145 of Winter 2013), but a negligible amount of dissolved molecular hydrogen. We estimate a lower bound of $\mbox{Si/O}\gtrsim 0.3$ based on Mid-Atlantic Ridge basalts (Winter, 2013). In the current study, we fix $\mbox{Si/O}=0.3$.
@@
Figure 7: Outgassing calculations informed by data-driven radius valley constraints. Left column: reduced conditions corresponding to $\log{f_{\rm O_{2}}}$ and $\log{f_{\rm S_{2}}}$ being uniformly sampled between 11-11 to 1-1 (cgs units; corresponding to $10^{-17}$ to $10^{-7}$ bar). Right column: oxidized conditions corresponding to $\log{f_{\rm O_{2}}}$ and $\log{f_{\rm S_{2}}}$ being uniformly sampled between 1-1 to $2$ (cgs units; corresponding to $10^{-7}$ to $10^{-4}$ bar). The top row shows the distributions of mean molecular weight, while the bottom row shows the corresponding oxygen fugacities relative to the iron-wüstite (IW) buffer (which is temperature- and pressure-dependent). The elemental abundance of carbon is assumed to be solar (see text for more details).
@@ 3.3 Representative case studies of hybrid atmospheres
where $G$ is Newton’s gravitational constant. We note that $10\mbox{ kbar}=1\mbox{ GPa}$. While seemingly large, such pressures are still much lower than what are required to render hydrogen metallic ($\sim 100$ GPa).
@@
There is a trend of the mean molecular weight increasing from $\mu\approx 2$ to $\mu\approx 16$ as the melt temperature decreases. This is slightly more pronounced for Earth-like, compared to the sub-Neptune-like, surface pressures. Nevertheless, the dominant effect is the diminished temperature, rather than the diminished pressure. Physically, these two properties are closely related: as the core becomes less massive, it retains less of its hydrogen envelope due to weaker gravity. This in turn allows the exoplanet to cool more quickly.